Further From Center has 11,400 shares of common stock outstanding at a price of $48 per share. It also has 275 shares of preferred stock outstanding at a price of $90 per share. There are 300 bonds outstanding that have a coupon rate of 6.7 percent paid semiannually. The bonds mature in 29 years, have a face value of $2,000, and sell at 108 percent of par. What is the capital structure weight of the preferred stock?

  • .0970

  • .4485

  • .5312

  • .0656

  • .0203
Answer #1

Value of Common Stock = 11,400(48) = $547,200

Value of Preferred Stock = 275(90) = $24,750

Value of Debt = 1.08(2,000)(300) = $648,000

Weight of Preferred Stock = 24,750/(547,200 + 24,750 + 648,000)

Weight of Preferred Stock = 0.0203